This is the "Renovation Updates" page of the "Renovation Updates - Maps and News" guide.
Alternate Page for Screenreader Users
Skip to Page Navigation
Skip to Page Content

Renovation Updates - Maps and News  

Last Updated: Jul 27, 2015 URL: Print Guide Email Alerts
Renovation Updates Print Page


The map reflects what the floor currently looks like.  


General Renovation Update

9/10/2012 -- Based on latest project update meetings

South-side book stacks: Except for some final shelves which need to be cut down in height, all shelving is complete. Facilities is working on these. When it's all done, Circulation will get collection into final configuration.

5th floor library entrance: UHD sticker put on one glass panel beside security gates, but it started to "bubble," so action on this stopped till problem can be resolved. Sign hung in outer lobby. Bulletin boards put beside entrance doors. (Acquiring more bulletin boards for inside library.) Billy will work with Facilities and IT to get one or two digital signage TV's put up to be viewable as people enter. Continuing to seek display cases. Panel to act as barrier on side of Security Desk in by end of September, then power and data will be added. UHD Presidential portraits will be hung on library south side this fall. Facilities to provide standard clock for outer lobby.

Public Space Furniture: Carrel design agreed on, will be ordered. Low "coffee" tables -- will seek a couple of more samples to look at. More 2-person tables to be ordered. Facilities still working on final layout of furniture, basing it on student use. Cannot decrease number of copiers due to contract, but will move "extras" to less-used areas. Facilities will remove "bubbles" from carpet in back hallway, move copiers and a little furniture into that area. Scanner and computer tables on order, still to come, will put them in final locations when they arrive. Christine working on signage additional to what Facilities does as a standard.

Rockwell Room and Coffee Shop: Finishes, carpet and furniture being finalized for Rockwell Room. Facilities presenting finishes and colors for Coffee Shop this week. Group to work on name for Coffee Shop to meet this week.

Circulation Desk: Cutting-down-desk project completed in August.

4th Floor: Chris, Billy, and Elena to work with Megan on layout and furnishing for emptied space in computer lab. Will try to combine workstations, study furniture, wireless printing (if approved by IT), and more scanners. Will buy a few more computer chairs for the lab. Will consider adding a little more furniture to Reading Room.




Shifting update:

1) The periodicals are about 50% moved to the South Side, we have temporarily run out of shifting space.  

2) J-PR is in it's permanent location on the North Side

3) A-HE is in it's permament location on the South Side

3) S-Z is in it's permanent location in the Northwest Corner.  We will continue this slower paced backwards shifting with the R's this week.  

4) We have begun moving Juvenile.  Juvenile Spanish is completely moved to the new low shelving.  After the rest Juvenile is finished - we will be the K-12 Textbooks in the additional low shleving by Juvenile. The Big Books (TBO, DT/Tesxtbook Oversized) will be rolled over sometime today. 


Moving the browsing periodicals today!  


Please send students to the Circulation desk for help finding items in the stacks.  

7/3/2012 - 10:42am

We have begun shifting the A's and B's to the Southwest corner.  Please let patrons know to ask at the circulation desk for assistance finding materials at this time.


5th floor doors are in working order.


The 5th floor doors entry way doors are working and locking from the inside but are not key accessable yet from the out side.  They are open to the public during working hours. (8-5).  If you plan to arrive before 8am, please continue to come in through the 4th floor until further notice. 

3/19/2012 - 10:25am - Done shifting

We are done shifting for the time being.

3/6/2012 - 11:18am - Shifting the R's

We are shifting in the R's today, expect to be finished by the end of the day.

2/22/2012 - 1:41pm - Shifting the P's

We started shifting the P's to the newly carpeted Southeast corner.  Shifting should be complete around Spring Break.

1/23/2012 - 12:09pm - New Books now available near Leisure Reading.  

The New Book shelves are now located near the A/V and Leisure Reading collections.

12/12/2011 - 11:21am - Old Circulation Desk demolition completed; New Books temporarily unavailable.  

Passage is available to the Administrative Offices and the Staff Break Room. The New Books should be temporarily relocated to the south side sometime this week.

11/21/2011 - 9:16am - Area connecting North and South towers on the west side is blocked off for abatement.  

Patrons must travel between towers on the east side by the circ desk.

11/7/2011 - 8:57am - Area on South Side Blocked Due to Abatement

The middle sector of 5th floor, south side is blocked off due to abatement, and will not be opened until Monday, November 14 (UPDATED DATE: WEDNESDAY, NOVEMBER 16). The restrooms are still accessible, but entrance to the southwest corner should be via the north side. Air conditioning is not operating in the south side public areas.

11/2/2011 - 3:58pm - Southside Catalog terminal relocated

The catalog terminal is relocated to the top of the staircase to the lab.  It is NOT operational yet.  I will let you know when it is. 

11/2/2011-3:56pm - Oversize DVD's on the shelf

We moved the oversized DVD's to media shelving with the rest of the DVD collection.  They are shelved separatly on the row directly behind the regular sized DVD's.    

10/17/2011 - 1:45pm - Moved through the N's

A-N is now located on the North Side.  

10/3/2011 10:58 am- DVD's and VHS merged and moved

The DVD's and VHS are merged together and moved to the open stacks as indicated on 9/30/2011.

10/3/2011 10:58 am - Special Collections Moved

The Special Collections have moved to new ILL Office

9/30/2011 - DVD's in the open stacks

The DVD's are moving to the open stacks today by the staircase to the 4th floor in front of leisure reading.

9/27/2011 3:08 - Copy Machines Moved

The copy machines by the elevator at the top of the staircase have been moved.  On is next to the cd cabinet and the others are next to the north side restrooms.

9/26/2011 8:50am - Beginning the HV move

We began moving the HV collection this weekend.  We will update the signs we we have completed the part of the move.

9/20/2011 9:08am - HE-HQ Relocated

These collections have been relocated to the North side.

8/29/2011 2:45 pm - HD - HE Relocated 

These collections have been relocated to the North side.

6/27/2011 11:16am - Juvenile and the K-12 Textbooks relocated

These collections have been relocated to the North side behind the periodicals.

6//15/2011 11:12 am - Copy Machines on the 5th Floor Moved

The copy machines on the 5th floor were moved to the South wall of Elevator 6 by the old North Circ desk (up the stairs, and to the left!)

6/15/2011 7:52am - Catalog in Juvenile removed

The catalog terminal in the Juvenile section was removed for construction.

6/03/2011 9:56am - Copy machine moved downstairs

 A second copy machine has moved downstairs.  It is the same model as the old one.

5/31/2011 10:48am - Special Collections Moved

The Special Collections have moved to S505E - the Circ/ILL office.

5/31/2011 10:02am- HC moved 

HC has moved to the new section as well.

5/26/2011 3:45pm - H, HA, HB moved

H, HA, and HB have moved to the new area behind the G's.  

5/24/2011 3:59pm - Northwest Study Corner under construction

The Northwest corner is under contruction.  The tables and chairs in this area have been removed.  

5/19/2011 9:55am 5th Floor Study Rooms demolished

The 2 study rooms located on the 5th floor were demolished this week.  

5/16/2011 10:44am CD Cabinet Moves

The CD Cabinet moved to behind the elevator next to the restrooms on the 5th floor.

5/16/2011 Construction Noise

Expect lots of construction noise for the next few weeks, since Facilities is taking advantage of the break when there are fewer students here.

4/29/2011 5:07pm - Catalog Terminals on the 5th Floor - Moved

The catalog computers on the 5th floor have moved closer to the south circ desk, right next to the New Book shelves.

4/26/2011 2:29pm - Periodicals Moving 

We will begin shifting periodicals tomorrow.  The will start with the A's behind the Current Periodicals by technical services.

4/26/2011 2:29pm- Circulation and ILL move complete

Interlibrary Loan, Brandon, Dashauna, and Michael are now working out of S505E.  Melissa and I are now working out of the old admin offices.  The North Circ desk is empty.  All operations are now out of the South Circ Desk.  

4/21/2011 ILL Move 

Interlibrary Loan is moved to S505E (the old instruction room).  We will get redirection signs up today.  

4/20/2011 10:29am - Paper Shredder

The paper shredder is now located in the supply closet in the old admin offices.

4/19/2011 3:39pm - Hallway to H!

The wall is now built around the Southeast corner of the 5th floor.  There is a Hallway built to so students may get to the H's.  I updated the Map to reflect todays changes.

4/19/2011 3:39pm- Circulation now at the south desk

The Circulation check out terminals are now located at the South desk.  The rest of circulation is in the process of moving and will hopefully be done tomorrow.

4/18/2011 10:02am - Circulation and ILL moving.  

We are moving to the old instruction room beginning today, and will hopefully be completed by Friday.  This includes moving the circ desk to the south side.  When each phase is completed, I will let you know.

4/18/2011 9:30am - Duct work on Tuesday

From Abraham "Fyi – we are having contractor install metal duct on Tuesday, April 19, 2011 in the area where the ceiling is open on the north side next to the copiers.  There will be minimal impact to the library." Update: Looks like they're doing this today, Thursday, April 21.

4/18/2011 9:30am - Abatement from this weekend completed

The abatement planned for this weekend in the southeast corner is completed. It was completed Sunday and the A/C was not turned on until today. However, the smell, still pretty strong, does not seem to affect anything except that immediate area.

4/14/2011 12:42pm - AV Equipment Relocated

The AV equipment has been relocated to the north side by the microfilm machines.  There is one study carrel with a 2 VHS players and 2 DVD players.

4/14/2011 7:52am - Status of the collection

I'll post something here anytime we make a notable change.  Currently the call numbers A-GV have been relocated in front of the circulation behind the leisure reading, vhs, audio, and oversize

4/13/2011 10:09am - Abatement scheduled for this weekend

Asbestos abatement is scheduled for this weekend in the Southeast corner of the 5th floor.  They will begin setting up around 5pm Friday, Saturday, and Sunday and work overnight, leaving each morning before we open.  

4/13/2011 10:05 am - AV Equipment Temporarily out of commision

The AV equipment along the South Side windows (vcr's, dvd players, cd players) is in the process of being relocated to the North Side by Tech services.  Due to low use, we are not putting the CD players back on the floor.  

4/13/2011 10:02am - Circulation and ILL on the move!

Circulation and ILL will begin moving offices when Ron returns next Tuesday.  Melissa and Christine will be in the old admin offices with Elena and Anne. Circulation and ILL staff will be in the old instruction room next to the old admin offices.  Immediately following the staff move, the Circulation Check Out desk will move from the desk on the North Side by the stairs to the South side while the area is renovated.  

4/13/2011 10:01am - Welcome!

Welcome to the Renovation LibGuide!  I will post some regular updates here, maps, and some contact info if you have questions :).


Loading  Loading...